A source of light O is located at a distance of 8cm from a plane mirror. The reflected ray O’  is detected 2 cm from the mirror at a vertical displacement of 12 cm. The point of reflection is at a distance of:

Grade:8

1 Answers

Arun
5 years ago
Dear Aditya
 
Sin i = 12-x/8Sin r =x/2Sin i = Sin r12-x/8 = x/2x = 12/5 = 2.4 cm from Y
 
Regards
Arun